A survey was conducted to measure the number of hours per week adults spend on home computers. In the survey, the number of hours were normally distributed, with a mean of 7 hours and a standard deviation of 1 hour. Find the probability that the hours spent on the home computer by the participant are less than 4.5 hours per week.
1
Expert's answer
2019-10-29T11:05:00-0400

Let "X=" the number of hours per week adults spend on home computers: "X\sim N(\mu, \sigma^2)"

Then


"Z={X-\mu \over \sigma}\sim N(0,1)"

Given that "\mu=7" hours and "\sigma=1" hour.


"P(X<4.5)=P(Z<{4.5-7 \over 1})=P(Z<-2.5)\approx0.0062"

The probability that the hours spent on the home computer by the participant are less than 4.5 hours per week is "0.0062" .